Suppose the money demand in an economy in which no interest is paid on money is M^d/p = 500+0.2Y-100i,


a) you're told that price =100, Y=100, and i=0.10.

Find the real money demand, nominal money demand and velocity.

b) the price doubles from p=10 to p=20. Find the real money demand, nominal money demand, and velocity.



1
Expert's answer
2021-05-30T14:18:40-0400

Solution:

Md = P * L (R, Y)

Where: P = Price level = 100

            L = Liquidity function

            R = Interest rate = 0.10

            Y = Real output = 100

Md = 500 + 0.2Y – 100i

a.). Real money demand:

Substitute the values in the demand function:

Md = 500 + 0.2(100) – 100(0.1)

Md = 500 + 20 – 10

Md = 510


Nominal money demand = Real money demand "\times" Price

Nominal money demand = 510 "\times" 100 = 51,000


Money velocity (MV) = Price (P) "\times" Real output (Y)

Money velocity (MV) = 100 "\times" 100 = 10,000


b.). An increase in P = 10 to P = 20:

Real money demand:

Substitute the values in the demand function:

Md = 500 + 0.2(100) – 100(0.1)

Md = 500 + 20 – 10

Md = 510


Nominal money demand = Real money demand "\times" Price

Nominal money demand = 510 "\times" 20 = 10,200


Money velocity (MV) = Price (P) "\times" Real output (Y)

Money velocity (MV) = 20 "\times" 100 = 20,000